OpenStudy (anonymous):

Write the slope-intercept form of the line that passes through the point (1, 0) and is parallel to x - y = 7. Help?

jimthompson5910 (jim_thompson5910):

Anything parallel to x - y = 7 will be in the form `x - y = C` where C is any constant to find the value of C, plug in (x,y) = (1,0) which is the given point x - y = C 1 - 0 = C 1 = C C = 1 so we have x - y = 1 after replacing C with 1 Now solve for y in x - y = 1